OpenStudy (anonymous):

Evaluate the expression if a = 3 , b = 6. and c = 1 2b / c _______ 1 - a @E.ali

OpenStudy (anonymous):

A. –6 B. 6 C. 24 D. –24

OpenStudy (anonymous):

i think its b i just wana be sure

OpenStudy (anonymous):

no i got it its A!!!

sam (.sam.):

Good!

OpenStudy (akashdeepdeb):

All you need to do here is JUST substitute the values given to you in the expression. \[a=3,b=6,c=1\] And we have to calculate \[\frac{ \frac{ 2b }{ c } }{ 1-a }\] OR \[\frac{ 2b }{ c(1-a) }\] Now put these values IN THE Expression. \[\frac{ 2*6 }{ 1(1-3) }\] Which gives us! \[\frac{ 12 }{ -2 }\] Understood this? :)

OpenStudy (anonymous):

Got it ???